(b) Find all the values of t E R for which T is positive definite. 2 -1 (i) T = -1 -1 t 2 ; 2 (ii) T = 3 + 4 0
Answer #1

NTT tot NTL 2 t t 2 NTT 2. t - 1 + 2 2 (4-6) +1(-2++) -1 (-t+2) 8-21-2+t+t-2 = -26ht2Ety so. The matsin is positive define ifowt ty o 4t Itlso 4 7 = t_970 tx 9 +>3; +4-3③ 0 TMO too and = t (t-16)-3(3+). t(€16-9) t(€ 25 0 ²,25 t> 5 t <-5 from ① and 2